OpenStudy (anonymous):

What is the slope of the line which passes through (–6, 0) and (4, 0)

OpenStudy (katrinakaif):

Use teh equation (y2-y1) / (x2-x1) as (-6,0) = (x1,y1) and (4,0) = (x2,y2)

OpenStudy (anonymous):

so the answer is 2?

OpenStudy (anonymous):

m=y2-y1/x2-x1

OpenStudy (anonymous):

Slope is defined as the change in y over the change in x. So you want to divide the difference between your y values and x values. So it would be (0-0)/4-(-6). Which would be 0.

OpenStudy (katrinakaif):

It would result as 10 /0 which equals 0

OpenStudy (anonymous):

oh I see it now thanks

OpenStudy (anonymous):

It would actually be 0/10. 10/0 would be undefined.

OpenStudy (unklerhaukus):

slope = rise over run =(0-0)/(4-(-6)) =0